Sandra lives in Anuradhapura and runs a business that sells mangos. In an average year, she receives Rs.3.6 million from sales of mangos. Of this sales revenue, she must pay the production and wholesale cost of Rs. 1,400,000. She also pays wages and utility bills totalling Rs. 1,100,000. If she does not operate this mango business, she can work in an accounting firm and receive a monthly salary of Rs. 70,000. She owns her showroom. If she chooses to rent it out, she will receive Rs. 5,000 in rent per month. Assume that the value of this showroom does not depreciate during the year. To commence the business, she has released Rs.1 million from her fixed deposit. The annual interest rate for her fixed deposit is 18%. No other costs are incurred in running this mango business. What is Sandra’s accounting profit (used for tax purposes)? What is her economic profit


1
Expert's answer
2021-09-04T15:17:38-0400

Accounting profit=total revenue -explicit cost

"=3600000-(1400000+1100000)\\

=3600000-2500000\\

=1100000"

Economic profit=Total revenue-explicit cost-implicit cost

"=3600000-(1400000+1100000)-(70000+5000)\\

=3600000-2500000-75000\\

=1025000"